OpenStudy (anonymous):

Tyler High School has a total of 101 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many girls play sports at this high school? a. 23 b. 31 c. 70 d. 80

OpenStudy (igreen):

If 23 girls: 23 \(\times\) 2 = 46 46 + 8 = 54 54 + 23 = 77. That's wrong. If 31 girls: 31 \(\times\) 2 = 62 62 + 8 = 70. 70 + 31 = 101. So the answer is 31.

OpenStudy (anonymous):

ok thanks!

OpenStudy (yanasidlinskiy):

By the way, It is a substitution method:D